# Representative Action Selection for Large Action Space Meta-Bandits

### v3

@@ 3.1 Regret Bounds of Algorithm
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof sketch. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, choose $a_{\ell}\in r_{\ell}\cap\mathcal{A}$ as the representative point. If $r_{\ell}\cap\mathcal{A}=\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$. The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset. Then, for each $\ell\leq m$, define a Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, we consider two cases. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$. If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is bounded by $\max_{a\in\mathcal{A}_{\text{full}}}\mu_{a}$. $\square$
The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 7 is maximized by the uniform $q$, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Supplementary. Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
The connection between Theorem 6 and 7 is insightful:
Remark 8.
Consider placing a grid over the action space and defining a partition $\mathcal{R}$ by assigning each point in the space to its nearest grid point. In this way, the grid acts as a reference subset with respect to the partition $\mathcal{R}$. The regret upper bound in Theorem 6 applies to this grid, as it holds for any reference subset. Meanwhile, Theorem 7 applies to any partition, including $\mathcal{R}$. As a result, we obtain a regret bound for Algorithm 1 that exceeds the grid’s bound by only one additional term:
@@
$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
where the part $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]$ is simply a constant, as the action space $\mathcal{A}_{\operatorname*{full}}$ is fixed.
If the distribution $q$ is highly concentrated, then the expectation $\mathbb{E}_{q}[(1-q(r))^{2K}]$ is small, making the extra term negligible. In this case, the benefit of constructing an explicit grid—often a non-trivial task when the action feature vectors are unknown or high-dimensional—is limited.
Generally speaking, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified. We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big\{m\!\in\!\mathbb{N}:\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big\}.$
Theorem 9.
Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 6, and another absolute constant $c>0$,
@@
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
The partition-independent upper bound in Theorem 9 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space. Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number. The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number. As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
The matching lower bound for Theorem 7, as well as the upper bound under the general sub-Gaussian process assumption for Algorithm 1, are provided in the Supplementary.
@@ 4 Generalization and Empirical Validation
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore treat the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$: $f(a):=\mu_{a},\forall a\in\mathcal{A}_{\operatorname*{full}}$. We show that our framework applies when the outcome functions lie in a reproducing kernel Hilbert space (RKHS) (williams2006gaussian, ). Consider a positive semidefinite kernel $k:\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$. By Mercer’s theorem, for a non-negative measure $\mathbb{P}$ over $\mathcal{A}_{\operatorname*{full}}$, if the kernel satisfies $\int_{\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}}k^{2}(a,a^{\prime})d\mathbb{P}(a)d\mathbb{P}(a^{\prime})<\infty,$ then it admits an eigenfunction expansion: $k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$ where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome functionff be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. The function $f$ can be rewritten as
$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
Assuming the outcome function is generated from a kernel inherently implies that actions are correlated. To study the effect of varying correlation levels among actions, we use the parametric RBF kernel to sample outcome functions from a RKHS: $k(a,a^{\prime}):=\exp\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right)$ , where $l$ is a length-scale parameter that control the dependence among actions. By (kanagawa2018gaussian, , Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$. Note that although the kernel is used to generate outcome functions, the kernel itself is not revealed; consequently, $a$ and $a^{\prime}$ are regarded merely as indices rather than as inputs to the kernel.
Figure 1: Comparison on solving equation 5, selecting $K=5$ distinct actions from 15 grid points over $[0,2]$, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales. Left: Expected regret over 50 repetitions of our method with exhaustive search (green), compared with Thompson Sampling (TS; yellow), Upper Confidence Interval (UCB; orange), and Successive Halving (SH, blue) applied to the set of super-arms. TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls. Right: SH’s expected regret (blue) and pull counts (yellow bars) per round, compared with our method (green). SH requires nearly 10,000 pulls—about three times the number of super-arms—to match our method.
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in equation 1. Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
@@
In general, the problem is non-trivial since the objective involves an expectation over random variables with an unspecified distribution, and the inner expression is a maximum over a subset of potentially correlated variables. Consequently, small changes in $\mathcal{A}$ can cause non-smooth variations in the maximum. Our method is computationally efficient, avoiding the combinatorial complexity.
We assess performance in identifying near-optimal solutions by selecting $K=5$ actions from an action space of 15 grid points in $[0,2]$. The configuration yields a moderate number of super arms, allowing approximation of the optimal subset using Successive Halving (SH), which, however, becomes impractical for larger action spaces. Outcome functions are generated with an RBF kernel at length-scales $l\in\{0.5,1,\dots,4\}$. We compare the proposed method against Thompson Sampling (TS) and Upper Confidence Bound (UCB). In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where exhaustive search is used over the action space to find the optimal action $a^{*}(\theta)$. For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$. Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective. In TS/UCB methods, we adopt a bandit feedback setting: at each round (corresponding to a bandit instance $\theta$), the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$). Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the prior set to $\mathcal{N}(0,1)$. In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in karnin2013almost , $N\log_{2}N\approx 37{,}000$) over a few rounds. In each round (corresponding to a bandit instance $\theta$), the number of remaining super arms is halved. This process continues until only the best-performing super arm remains or the budget is exhausted. We evaluate expected regret of all methods over $10^{5}$ additional randomly sampled outcome functions.
The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown, the expected regret decreases as the length-scale increases. The length-scale $l$ controls the effective number of approximately independent actions: when $a$ and $a^{\prime}$ are far relative to $l$, $k(a,a^{\prime})$ is negligible, so $\mu_{a}$ and $\mu_{a^{\prime}}$ are nearly uncorrelated. In the limit, $\{\mu_{a}\}_{a\in\mathcal{A}_{\text{full}}}$ forms a collection of i.i.d. zero-mean, unit-variance variables.
The SH in left subplot of Figure 1 is used to approximate the ground-truth optimal subset and our method does give similar performance as SH. The right subplot highlights SH’s impracticality: we ran our method (green) and SH (blue) for 50 repetitions with fixed $l=1$, tracking the current best super arm and total super arm pulls after each SH round. The blue curves and shades represent the mean $\pm$ one standard deviation of SH’ expected regret, while yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round. Green curves show our method’s expected regret. SH matches our performance after three rounds but requires 9005 pulls—roughly three times the number of super arms—making it intractable for large $\mathcal{A}_{\operatorname*{full}}$.
@@
Figure 2: Comparison of our method (green), CTS (brown), and CUCB (red) on solving equation 5 with $K=10$ actions chosen from 500 grid points over [5,5][-5,5], with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales. In our method, EpsilonNet+TS, the “argmax” step in Algorithm 1 is replaced by TS run for 300 rounds. CTS is run for 3000 rounds and CUCB for 2000 rounds. All methods use prior variance 11. Left: expected regret over 30 repetitions. Right: runtime (purple, left axis) and number of arm pulls (gray, right axis) over 30 repetitions.
Further, we demonstrate the generality of our method (denoted EpsilonNet+TS), where the “argmax” step in Algorithm 1 is replaced by TS. We consider the same setup as in Figure 1, but with $K=10$ actions chosen from 500 grid points over [5,5][-5,5]. In our method, for each bandit instance $\theta$, we run TS for 300 rounds, selecting an action $a$ and observing $\mu_{a}$ in each round; the final action is taken as an approximation of the optimal action $a^{*}(\theta)$. We repeat this process until $K=10$ distinct actions are identified. We compare against two heuristics: combinatorial TS (CTS) and combinatorial UCB (CUCB). CTS is run for 3000 rounds to match the runtime of our method, while CUCB is run for 2000 rounds due to its slower execution and additional initialization step. Both CTS and CUCB operate in the semi-bandit feedback setting: in each round (corresponding to one bandit instance $\theta$), an outcome function is sampled and $K=10$ actions $\mathcal{A}$ are selected, with the rewards $\mu_{a}(\theta)$ revealed for all $a\in\mathcal{A}$. The prior for TS, CTS, and CUCB is set to $\mathcal{N}(0,1)$, consistent with the ground truth.
In Figure 2, the left subplot reports the expected regret and its standard deviation for EpsilonNet+TS (green), CTS (brown), and CUCB (red), averaged over 30 repetitions. The poor performance of CTS and CUCB is expected, as noted in related works. The variance of EpsilonNet+TS could be further reduced by running more TS rounds or using a more accurate action-identification method, since we currently use a Gaussian prior—modeling each action’s expected outcome $\mu_{a}$ as an independent Gaussian random variable—rather than a Gaussian process prior, in order to accommodate the most general case where $\mathcal{A}_{\operatorname*{full}}$ is an index set. In the right subplot, we fix the length-scale $l=1$ and evaluate runtime and arm pulls of the three methods over 30 repetitions. The left axis (purple) shows average runtime $\pm$ one standard deviation, and the right axis (gray) shows the corresponding number of arm pulls. EpsilonNet+TS requires 3000 pulls, since it selects $K=10$ optimal actions, with each action identified by running TS for 300 rounds. CTS has similar runtime but needs far more pulls, as it runs 3000 rounds with $K=10$ actions pulled per round.
Conclusion and Future Work
We proposed a framework for selecting a subset of correlated actions. A simple algorithm was introduced and shown to effectively identify near-optimal subsets. Future work could develop a stopping criterion. When subset size is flexible but sampling new bandit instances is costly, ideas from species discovery (roswell2021conceptual, ) may be useful—treating the discovery of a new action as analogous to discovering a new species.
Appendix Summary: The appendix is divided into four parts.

### v4

@@ 4.2 Regret Bounds for the Algorithm
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\left(\mu_{a}-\mu_{a_{\ell}}\right)\right]+C\epsilon\sqrt{\log m}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},\end{split}$
where $C>0$ is an absolute constant.
Note that the expectation $\mathbb{E}_{\theta,\mathcal{A}}$ in this result is taken with respect to both the algorithm’s randomness (i.e. the sampled set $\mathcal{A}$) and the distribution over $\theta$.
In the following remarks we discuss the different error terms appearing in Theorem 4.5.
Remark 4.6 (The Cluster Complexity Term).
The term $\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]$ in Theorem 4.5 measures the size of the cluster, and is also known as the Gaussian width of $r_{\ell}$ (see (Vershynin, 2018), Chapter 7). There are two standard ways in which we can bound this term: First, for any finite $r_{\ell}$ we have
@@
$\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]\leq\frac{\operatorname{diam}(r_{\ell})}{2}\sqrt{\log|r_{\ell}|}\leq\frac{\epsilon}{2}\sqrt{\log\left|\mathcal{A}_{\operatorname*{full}}\right|},$ (8)
thus yielding a scaling with the diameter and a square-root-logarithmic dependence on the size of the actions set. See Supplementary K additional details on (8). In addition, we have the bound
$\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]\leq\operatorname{diam}(r_{\ell})\frac{\sqrt{n}}{2},$ (9)
where $n$ is the ambient dimension of the action set, i.e. $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$ (see (Vershynin, 2018), Chapter 7). This bound is independent of the cardinality of the action set and may be applied if $n$ is known to be small. However, as discussed in Section 3.1, we do not assume that $n$ is small. We use both bounds in Theorem 4.9 below.
Remark 4.7 (Sampling Correction Term).
The sampling correction term $\mathbb{E}_{q}\big[(1-q(r))^{2K}\big]$ in Theorem 4.5 decays exponentially with $K$. Here we analyze this decay in more detail. Write
$\mathbb{E}_{q}\big[(1-q(r))^{2K}\big]$ $=\sum_{r}q(r)(1-q(r))^{2K}$
$\leq\sum_{r}q(r)e^{-2Kq(r)},$
where we have used $1-x\leq e^{-x}$ for $x\in[0,1]$. For each cluster $r\in\mathcal{R}$, the summand $q(r)e^{-2Kq(r)}$ is small when $q(r)$ is small and decays rapidly when $q(r)$ is large. For any threshold $\tau\in(0,1)$, we split the sum into contributions from “large” and “small” probabilities:
$\sum_{r}q(r)e^{-2Kq(r)}\leq e^{-2K\tau}+\sum_{r:q(r)\leq\tau}q(r).$
Thus, to ensure $\mathbb{E}_{q}[(1-q(r))^{2K}]\leq\varepsilon$, it suffices to choose $\tau$ such that $\sum_{r:q(r)\leq\tau}q(r)\leq\varepsilon/2$ and set
$K\geq\frac{1}{2\tau}\log\!\left(\frac{2}{\varepsilon}\right).$
Remark 4.8 (Linear Bandits and Known Gaussian Process Regret Bounds).
As discussed in Section 3.1, when dimension $n$ is fixed, we can view the problem setup (5) as that of Bayesian linear bandits. In Section 1.1 we have mentioned two motivations for the action subset selections problem: The intrinsic interest in reducing the size of the action set for purposes such inventory maintenance, and a possibility of faster learning in a new bandit instances, when the action set is reduced from $\mathcal{A}_{\operatorname*{full}}$ to $\mathcal{A}$, see (2). While the first scenario holds for linear bandits as well, the second scenario may be weaker, since in linear bandits the regret may be bounded either as $O(\sqrt{n\cdot\log\left|\mathcal{A}_{\operatorname*{full}}\right|}\sqrt{T})$, (Russo & Van Roy, 2016), or as $O(n\sqrt{T})$, (Dani et al., 2008). Thus, in these situations the regret depends at most only logarithmically on the size of $\left|\mathcal{A}_{\operatorname*{full}}\right|$, if the dimension $n$ is known to be small. However, as discussed earlier, we do not require the dimension to be small, and our bounds depend on the actual geometry of the set instead of relying on global constraints such as the dimension. Thus our bounds may remain small even when $n$ is arbitrarily large.
Similar considerations also apply more generally to existing regret bounds for Bayesian bandits with Gaussian Process reward distributions (a problem also known as Bayesian Optimization), (Srinivas et al., 2009; Garnett, 2023). While these bounds are applicable with infinite $\mathcal{A}_{\operatorname*{full}}$, they involve global parameters of the model and assumptions on the kernel. For instance, for linear kernels the bound is $O(n\sqrt{T})$, while for RBF kernels one has $O(\sqrt{T\left(\log T\right)^{n+1}})$ - exponential in the underlying dimension, (Srinivas et al., 2009). Thus such bounds do not reflect the underlying clustering or non-linear low dimensionality of the data.
Even for small $n$, each round incurs non-trivial optimization costs. Both UCB and TS require solving an inner maximization over $\mathcal{A}_{\text{full}}$, whose difficulty depends on the action space structure. In linear bandits, UCB solves $\max_{a\in\mathcal{A}_{\text{full}}}\max_{\theta\in\mathcal{C}_{t}}\langle\theta,a\rangle$, which is NP-hard in general—even for convex $\mathcal{A}_{\text{full}}$ (Zhang et al., 2025)—but trivial for small finite sets. TS similarly requires $\max_{a\in\mathcal{A}_{\text{full}}}\langle\tilde{\theta}_{t},a\rangle$; while linear in $a$, this is costly for large finite sets, though near-sublinear costs $\tilde{O}(|\mathcal{A}_{\text{full}}|^{1-o(1)})$ are achievable under structural assumptions (Jun et al., 2017; Yang et al., 2022). For GP bandits, the inner problem maximizes a non-concave acquisition function (e.g., GP-UCB), which is generally intractable for continuous domains. With finite $\mathcal{A}$ of size $|\mathcal{A}|$, the per-step cost is $\tilde{O}(|\mathcal{A}|h^{2}+h^{3})$ for batch size $h$ (Calandriello et al., 2022). Continuous domains typically require discretizations for approximate maximization (Chowdhury & Gopalan, 2017; Salgia et al., 2021).
Figure 2: Comparison on solving (11) by selecting $K=5$ actions from 15 grid points in $[0,2]$, with outcome functions $f(a)=\mu_{a}$ sampled via an RBF kernel at varying length-scales. Left: Expected regret over 50 repetitions: our method with exhaustive search (green) vs. super-arm TS (yellow), UCB (orange), and Successive Halving (SH; blue). TS/UCB run for 3,000 rounds; SH uses a budget of 37,000 pulls. Right: SH’s expected regret (blue) and cumulative super-arm pulls (yellow bars) per round vs. our method (green). SH requires nearly 10,000 pulls—about three times the number of super-arms—to match our performance.
Finally, as discussed earlier, while Theorem 4.5 is stated for a fixed partition, for any $\epsilon>0$ we can construct a partition derived from an $\epsilon$-net, and to obtain a corresponding regret bound. We carry this construction out in the following result. Recall the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$ from (7).
Theorem 4.9.
Set $M:=\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})$ and fix $\epsilon>0$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for
$K\geq c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon),$ (10)
we have
@@
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;$ $C\epsilon\min\{\sqrt{n},\sqrt{\log|\mathcal{A}_{\operatorname*{full}}|}\}$
$+C^{\prime}\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
where $C,C^{\prime},c>0$ are absolute constants.
In particular, by taking $\epsilon\to 0^{+}$, and taking $K$ to satisfy (10), we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$.
A matching lower bound and an extension to the sub-Gaussian case are provided in Supplementary B, along with the full proofs of the results in this section.
@@ 5 Generalization and Empirical Validation
We now treat the expected rewards $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as evaluations of an unknown outcome function $f:\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$ over the abstract action set, where $f(a):=\mu_{a}$. Although our framework is general, we instantiate it in the experiments using outcome functions from an RKHS (Williams & Rasmussen, 2006), a function class for which our method is applicable and theoretically sound. The synthetic outcome functions are sampled from a Gaussian process with an RBF kernel (with length-scale parameter $\ell$), yielding outcome functions consistent with an underlying RKHS (Kanagawa et al., 2018). The action space corresponds to the input domain of the RBF kernel, while the kernel itself is not revealed to the learner.
Details justifying the applicability of our framework to RKHS, as well as the data generation procedure, are provided in Supplementary C.
Returning to the objective stated in the Introduction, we seek a subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ of cardinality $K$ that minimizes the expected regret defined in (1). Since $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is constant with respect to $\mathcal{A}$, this is equivalent to solving:
@@
This problem is generally non-trivial: the objective involves an expectation over random variables with an unspecified (and potentially correlated) distribution, and the inner $\max$ operator induces non-smooth dependence on $\mathcal{A}$. Nevertheless, our method efficiently circumvents the combinatorial complexity inherent in (11).
@@
Figure 3: Comparison of EpsilonNet+TS (green), CTS (brown), and CUCB (red) on solving (11) with $K=10$ actions from 500 grid points in [5,5][-5,5], using outcome functions $f(a)=\mu_{a}$ sampled via an RBF kernel at varying length-scales. In EpsilonNet+TS, the argmax oracle in Algorithm 1 is replaced by TS; CTS and CUCB run for 3,000 rounds. All methods use a $\mathcal{N}(0,1)$ prior. Left: expected regret (30 repetitions). Right: runtime (purple, left axis) and arm pulls (gray, right axis).
5.1 Comparisons at the Super-Arm Level
We compare our method against TS and UCB in terms of super-arm identification performance: we aim to identify near-optimal subsets of (11) by selecting $K=5$ actions from an action space consisting of 15 grid points in $[0,2]$. Each 5-action tuple is treated as a super-arm. This yields $\binom{15}{5}=3003$ super-arms—a moderate number that allows us to approximate the optimal super-arm using Successive Halving (SH), although SH becomes computationally infeasible for larger action spaces. Outcome functions are sampled from via an RBF kernel at length scales $l\in\{0.5,1,\dots,4\}$.
In our approach, we run Algorithm 1 until $K=5$ distinct actions are selected, using exhaustive search over the action space to compute the optimal action $a^{*}(\theta)$ for each bandit instance $\theta$. TS and UCB operate in a bandit feedback setting: at each round (i.e., for a new $\theta$), a super-arm $\mathcal{A}$ is selected, the payoff $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ is observed, and the policy is updated. We run both methods for 3,000 rounds—matching the total number of super-arms ($\binom{15}{5}=3003$). The payoff of a super-arm $\mathcal{A}$ is defined as $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ to align with the objective in (11); thus, these methods directly optimize $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$. Since this payoff is unbounded, we use a $\mathcal{N}(0,1)$ prior for both TS and UCB. For SH, we allocate a fixed budget of $N\log_{2}N\approx 37{,}000$ arm pulls (following (Karnin et al., 2013)), where $N=3003$ is the number of super-arms. In each round (corresponding to a new $\theta$), the set of surviving super-arms is halved based on empirical payoffs, continuing until one remains or the budget is exhausted. Finally, we estimate the expected regret of all methods over $10^{5}$ additional sampled outcome functions.
The left subplot of Figure 2 shows the expected regret of our method (green), TS (yellow), UCB (orange), and SH (blue). Solid curves and shaded regions denote the mean $\pm$ one standard deviation over $50$ repetitions. As observed, the expected regret decreases with increasing length-scale $l$. The length-scale $l$ controls the effective number of independent actions: for $|a-a^{\prime}|\gg l$, $\mu_{a}$ and $\mu_{a^{\prime}}$ become nearly uncorrelated. In the limit of large $l$, the outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\text{full}}}$ behave like i.i.d. $\mathcal{N}(0,1)$ variables.
The right subplot of Figure 2 highlights SH’s impracticality. We ran both our method (green) and SH (blue) for 50 repetitions with fixed $l=1$, tracking the best super-arm found and total super-arm pulls after each SH round. The blue curves (with error shades) show SH’s expected regret (mean $\pm$ std.), while the yellow bars (right y-axis) indicate cumulative super-arm pulls per round. SH matches our performance only after three rounds—but at roughly three times the total number of super-arms.
5.2 Oracle-Free Comparisons with CTS and CUCB
While Algorithm 1 relies on an oracle argmax step, this is often infeasible in practice. We replace it with a simple TS rule, yielding a fully implementable, oracle-free variant (EpsilonNet+TS). We compare this method against combinatorial TS (CTS) and combinatorial UCB (CUCB)—two practical approaches for choosing an action subset with $|\mathcal{A}|=K$—under the same setup as in Section 5.1, but now selecting $K=10$ actions from an action space consisting of 500 grid points over [5,5][-5,5].
Our approach runs Algorithm 1 until 10 distinct actions are identified. The oracle argmax step is approximated by running TS for 300 rounds; the action chosen in the final round serves as an approximation of $a^{*}(\theta)$. CTS and CUCB operate in a semi-bandit setting for 3000 rounds. In each round $t$, an instance $\theta_{t}$ is sampled, an action subset $\mathcal{A}_{t}$ of cardinality $K=10$ is selected, and rewards $\mu_{a}(\theta_{t})$ are observed for all $a\in\mathcal{A}_{t}$. The subset selected in the final round is taken as the algorithm’s output. All methods use a $\mathcal{N}(0,1)$ prior, consistent with the ground truth.
The left subplot of Figure 3, the left subplot shows the expected regret (mean $\pm$ standard deviation over 30 repetitions) for EpsilonNet+TS (green), CTS (brown), and CUCB (red). As observed, the poor performance of CTS and CUCB aligns with discussions in related work. The right subplot reports runtime and arm pulls across 30 repetitions with fixed $l=1$: the left axis (purple) shows average runtime $\pm$ one standard deviation, while the right axis (gray) shows total base arm pulls. EpsilonNet+TS uses exactly 3,000 arm pulls (300 TS rounds each $\times$ 10 base arms). CTS incurs comparable runtime but requires far more pulls (3,000 CTS rounds $\times$ 10 base arm pulls). CUCB is significantly slower and incurs additional pulls due to its mandatory initialization.
6 Conclusions and Future Work
We proposed a framework for selecting a small representative subset of actions from a very large shared action space in a family of bandit problems. Our simple subset-selection algorithm exploits reward correlations to identify representative actions with controllable performance loss. Natural directions for future work include developing data-dependent stopping rules for choosing the subset size K, and extending the setting from bandits to Markov decision processes to capture sequential decision making.
Supplementary Summary: The appendix is divided into four parts.